alexgame:

Meg is 6 years older than Victor. Meg's age is 2 years less than five times Victor's age. The equations below model the relationship between Meg's age (m) and Victor's age (v): m = v + 6 m = 5v − 2 Which is a possible correct method to find Meg's and Victor's ages? Solve m + 6 = 5m − 2 to find the value of m. Write the points where the graphs of the equations intersect the x axis. Solve v + 6 = 5v − 2 to find the value of v. Write the points where the graphs of the equations intersect the y axis.

alexgame:

hello again

SmokeyBrown:

Hello hello! So, you have two equations that both equal m, right? m = v + 6 m = 5v − 2 Imagine putting the m in the middle, like v + 6 = m = 5v - 2 Both sides are equal to each other, so this is just like saying v + 6 = 5v - 2 Then you could solve for v... I even think one of the answer choices says that too!

alexgame:

so its going to be c